Circular No. 8551

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005cq M. Riello, Osservatorio Astronomico di Padova (OAP); S. Valenti, Uni-

versitá di Ferrara; M. Turatto and S. Benetti, OAP; M. T. Botticella and E. Cappellaro, Osservatorio Astronomico di Capodimonte; and F. Patat, European Southern Observatory (ESO), report the discovery of a supernova (R = 21.9) on a CCD frame taken during a search performed with the Max-Planck-Institute 2.2-m telescope (+ Wide Field Imager) at the ESO on May 13. SN 2005cq is located at α = 9^{h}52^{m}00^{s}.48, δ = −20^{o}43′26′′.5 (equinox 2000.0); the supernova and the host galaxy are not visible on reference images taken with the same equipment in Feb. 2004 (limiting mag R ∼ 24). A spectrum (range 600–1100 nm, resolution 0.6 nm), obtained on May 23.46 UT with the ESO 8.2-m ANTU telescope (+ FORS2), shows 2005cq to be a type-Ia supernova ∼ 1 week past maximum light, when compared with the spectrum of SN 1994D (Patat et al. 1996, MNRAS 278,

111) and a redshift of 0.31 is assumed for the parent galaxy.

COMETS C/2005 J11, J12, and K4–K8 (SOHO) Following are the initial available positions for additional presumed

comets found on SOHO website images (continuation to IAUC 8549). All were Kreutz sungrazers except for C/2005 K4 (Meyer group). K. Battams writes that C/2005 J11 was diffuse with no tail, reaching mag ∼ 7.2 at 5.8R_{⊙} on May 15.243 UT. C/2005 J12 was tiny with no tail (too faint for photometry). C/2005 K4 was also tiny and barely above the background, and it showed no elongation. C/2005 K5 was very small, ‘teardrop’-shaped, and diffuse, reaching mag ∼ 8.2 at 5.9R_{⊙} on May 18.701. C/2005 K6 was very diffuse with no tail and barely above the sky background. C/2005 K7 was an extremely faint and diffuse companion to C/2005 K8, which in turn was diffuse and had a hint of a tail (reaching mag ∼ 8.2 at 5.6R_{⊙} on May 20.535).

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 J11 May 15.213 3^{h}32.^{m}2 +17^{o}26′ C2 BZ 2005-M39
C/2005 J12  15.910 3 34.9 +17 34 C2 TH 2005-M39
C/2005 K4  16.226 3 35.9 +20 21 C2 RK 2005-M39
C/2005 K5  18.685 3 45.3 +18 10 C2 HS 2005-M39
C/2005 K6  20.129 3 50.4 +18 24 C2 TH 2005-M39
C/2005 K7  20.504 3 52.4 +18 33 C2 KB 2005-M39
C/2005 K8  20.518 3 52.0 +18 35 C2 HS 2005-M39
2005 June 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green